# BLACKOUT: Data-Oblivious Computation with Blinded Capabilities

### v3

@@ Abstract.
[…]
=2ex
@@ 1. Introduction
Weak memory-safety and side-channel leakage are two significant security challenges for security-critical software, such as cryptographic libraries. Lack of memory-safety in programming languages such as C and C++, is one of the oldest, most persistent problems in computer security. The urgency of addressing the impact of memory-unsafe code at scale has grown under recent regulatory scrutiny (ONCD, 2024) leading cybersecurity authorities like US CISA (CISA, 2023) and UK NCSC (NCSC, 2024) to advocate for memory-safe languages like Rust (Klabnik and Nichols, 2018), and memory-safe hardware, such as Capability Hardware Enhanced RISC Instructions (CHERI) (Watson et al., 2019) as potential ways to eliminate entire classes of memory-safety vulnerabilities. Rust is poised to benefit from increased uptake thanks to the advocacy from regulators, but has not yet been proven in many industry sectors, and a complete rewrite of the vast body of existing C/C++ code in use today at a grand scale is economically impossible.
@@
Memory-safe hardware, such as CHERI, provides an alternative way to ensure memory-safety properties, particularly in systems where C and C++ will remain the dominant languages for the foreseeable future. CHERI effectively addresses memory safety, but does not inherently protect against side-channel attacks. Previous work to harden CHERI against transient execution attacks (Fuchs et al., 2024) does not protect against non-transient side-channel attacks. On the other hand, inherent resistance against side channels would also be effective against transient execution attacks that employ such side channels for inferring their effects on microarchitectural properties like cache state. In principle, CHERI-capable hardware could be complemented with existing data-oblivious computation solutions like OISA (Yu et al., 2019) or BliMe (ElAtali et al., 2024b) that protect against side channels. However, these approaches introduce considerable performance overhead because they require the addition of hardware-managed tag bits to memory. The adoption of hardware-accelerated memory tagging is hampered by significant meta-data storage and memory traffic overheads, since tags must be managed for every piece of data in memory.
@@
We plan to open-source the artifacts for the BLACKOUT complete hardware and software stack to enable reproducibility of our results and to facilitate further research by the community.
@@ 2.1. The CHERI capability architecture
{bytefield}
[…]
@@
CheriBSD is a modified version of the open-source FreeBSD operating system, designed to support CHERI-RISC-V both in emulation and on hardware. It is a fully functional OS prototype, demonstrating how CHERI support can be integrated into a conventional OS design. The CheriBSD kernel and userspace can be built in “pure-capabilityCHERI C/C++ which means all conventional memory pointers are replaced by corresponding capabilities by the CHERI-LLVM compiler. Cornucopia is integrated into CheriBSD via a shim layer, known as the malloc revocation shim (MRS), which sits on top of the underlying BSD libcs memory allocator. Consequently, CheriBSD enforces both spatial and temporal memory safety for CHERI-enabled software.
$2.$ Side-Channel Leakage
@@
A prominent example of a timing side channel is cache timing, which leaks information by measuring cache access latency. By comparing the latency to a predetermined threshold, software can determine whether an address is already cached by the CPU. Prime+Probe (Osvik et al., 2006), FLUSH+RELOAD (Yarom and Falkner, 2014) and Flush+Flush (Gruss et al., 2016) are example side-channel attacks using different techniques to make this timing difference dependent on secret data and extract it from the system. Transient attacks like Spectre (Kocher et al., 2019) and Meltdown (Lipp et al., 2018), expose new ways to access (architecturally inaccessible) secrets and leak them using side channels like cache timing.
Preventing side-channel leakage is difficult for software developers. Software is often written in high-level languages that abstract away the individual CPU instructions operating on secret data, leaving decisions to the underlying compilers and software stack. Even lower-level languages like C do not provide developers with sufficient controls to prevent side-channel leakage, as the compiler can introduce side channels in the generated assembly. Furthermore, even with seemingly side-channel-free assembly, hardware optimizations transparent to software (and often undocumented), such as speculation and out-of-order execution, can inadvertently leak secrets, as described above with the Spectre and Meltdown vulnerabilities. The challenge of preventing side-channel leakage is therefore present at every level of the software and hardware stack, and must be solved through hardware-software co-design.
@@ 2.3. Data-Oblivious Instruction Set Architectures
Data-oblivious computation refers to algorithms and software that process data in a manner where their control-flow and memory access patterns are not affected by their input data. Constant-time code, which hardens software against side-channels, relies on data-oblivious algorithms and that the CPU cycles consumed by individual hardware instructions are independent of their operands’ values111“Constant-time code” is a misnomer, as even side-channel-resistant code may still exhibit variable latency if the latency differences are not due to the (secret) data values. Common sources of such latencies include instruction-level latencies due to frequency scaling, or unpredictable contention of data buses for loads and stores.. Historically, developers of security-critical code sensitive to side-channels, e.g., cryptographic implementations, have relied on obscure information on instruction timing differences obtained through extensive experiments on different ISAs (Pornin, 2018a). In recent years, major processor manufacturers have began to document instructions with data operand independent timing (Intel, 2023a) and incorporate data operand independent timing modes into their designs (Intel, 2023b).
The challenge of ensuring the correctness of data-oblivious software implementations, however, remains. Verifying whether a program written in a high-level language is data-oblivious has a number of challenges (§ 9). Chief among them is that data-obliviousness can only be defined at a machine-code level. This observation has motivated the creation of data-oblivious ISAs that, similar to data operand independent timing modes, provide modes of operation for CPUs to ensure that instructions operating on sensitive data do so in a data-oblivious manner. Previous works use hardware-enforced taint tracking using “blindedness tags” (ElAtali et al., 2024a) and dedicated memory partitions (Yu et al., 2019). These approaches result in best-case performance overheads from 8% to 35% (worst case involving an order of magnitude slowdown). Memory tagging exhibits poor performance as the number of tag bits increase. Schemes such as the Arm architecture’s Memory Tagging Extension (MTE), used for memory-safety sanitation, limit the number of tag bits to 4. MTE further mitigates performance impact by allowing tag checks to occur asynchronously. But asynchronous tag checks are not practical for maintaining data-obliviousness as data-oblivious ISAs must be able to prevent data accesses before confidentiality is violated.
@@ 3. System and adversary model
We assume the same system model as in CHERI. We trust the OS to prevent memory containing sensitive information from being exposed to other processes after a process exits, either normally or as a result of a CHERI exception. We assume the memory-safety properties provided by CHERI, which prevent adversaries from tampering with the program’s control flow or directly inferring the contents of memory beyond that which is exposed in the course of normal program operation.
We also assume the same adversary model as in CHERI. In addition, we assume the adversary has the ability to observe side channels during a program’s execution, possibly from another process running on the system simultaneously. Timing side channels, such as cache timing and instruction timing, are therefore in scope. However, attacks that require physical access to the system, e.g., to measure power consumption, are out of scope. Attacks through direct memory access (DMA) peripherals are also out of scope.
@@ 4.1. Goals and Primary Challenges
To address these challenges, we propose enhancing the existing CHERI capability model with blinded capabilities which ensure data accessed through them is operated on exclusively in a data-oblivious manner. Memory managed by blinded capabilities is referred to as blinded memory. This avoids the need to propagate blindedness tags to memory thus overcoming the drawbacks of simply integrating an existing data-oblivious ISA with CHERI (ElAtali et al., 2024b; Yu et al., 2019).
@@ 5.1. Software Architecture
Figure 2 shows a high-level overview of the blinded capability software stack. Developers can indicate to the compiler that a variable is sensitive by annotating its declaration with the blinded attribute ➀. We refer to such variables as blinded variables. The compiler will then ensure that any accesses to blinded variables are permitted only through blinded capabilities, guaranteeing that the values of these variables will always be blinded when loaded into registers. We refer to registers holding blinded data as blinded registers. Blinded memory can be dynamically allocated using the blinded allocator application programming interface (API) ➁ which returns a blinded capability. Similarly, the compiler enforces that references to blinded memory cannot be assigned to non-blinded variables. Inside the compiler, we modify the Clang front-end to emit blinded attributes to the LLVM intermediate representation (IR) of the compiled program ➃. These are used by the blinded capability analysis ➄ and instrumentation passes ➅ to produce a blinded capability-instrumented application binary.
At run-time, the blinded-capability-enhanced application ➆ requires platform support in the form of the blinded allocator ➇ and an OS with a CHERI-capable kernel that has been altered to ensure its internal handling of capabilities does not trigger blinded capability violations. Other system software which does not make use of blinded capabilities does not require modification (apart from their adaptation to CHERI). We discuss the compiler enhancements and operating system changes in detail in § 6.2.
1#define __blinded [[clang::annotate_type("blinded")]] 2 3 __attribute__ ((blinded)) 4int data_oblivious_select(bool cond, int x, int y) { 5 6 bool __blinded c; // c declared blinded and 7 // accessed via blinded capability 8 9 int res; // res not declared blinded 10 // but blindedness is inferred 11 12 13 c = cond; // Uses store via blinded capability 14 // (argument already in register) 15 16 17 res = (a * c) + (b * (!c)); // HW propagates 18 // blindedness to res 19 20 21 return res; 22} Listing 1: Data oblivious conditional select function which returns one of the arguments x or y based on the value of cond demonstrating the inference capabilities of the blinded capability-enhanced compiler. ❸ Compiler infers res blinded from ❺❹ Compiler knows c is already blinded based on declaration❺ Compiler infers res is blinded from this assignment
@@ 5.2. Hardware architecture
Figure 3 shows the high-level overview of the hardware changes needed to support blinded capabilities (shown in dark gray) when applied to an out-of-order RISC-V core, such as MIT’s RISCY-$O$ (Zhang, 2025) or Bluespec’s Toooba (Nikhil, 2025) processor intellectual property (IP). Apart from the changes introduced by CHERI (shown in light gray) the majority of the changes necessary in the CPU are to facilitate in-CPU taint tracking. In the register file, the general-purpose capability registers are extended to hold an additional blindedness bit ➀ which tracks whether the register is blinded. The taint-propagation hardware logic ➁ will in turn ensure that any data derived from these blinded registers, e.g., through arithmetic operations, will also “be blinded”, i.e., it sets the blindedness bit of the destination register.
All blinded data is bound by a set of invariants (I1, I2, I3, I4 and I5), stemming from the need to enforce exclusive access to blinded memory (§ 4.2), and side-channel protection.
@@
I1 is enforced by hardware through additional blinded store and load checks ➂ in the memory execution pipeline which is responsible for the execution of any load and store instructions, whereas I4 and I5 are enforced by side-channel prevention logic integrated into any pipeline which is responsible for instructions with blinded operands ➃. Any attempt by software to violate I1, I4 and I5 will result in a hardware fault.
@@
In practice, developers usually do not wish to keep all of their data blinded indefinitely as some result of data-oblivious computation must eventually be extracted. One way to reveal the final, non-secret result of blinded computation is to relax invariant I3 in a controlled way to unblind the result. For example, since I3 is (in part) enforced by the blinded allocator, the allocator can expose a separate API for obtaining blinded results that, in contrast to, regular blinded memory returns two capabilities for a dedicated area for blinded results: one which is blinded, and used to write to the area, and the other non-blinded, but only usable for reading the result. However, the developer must use the blinded capability to the result area carefully to prevent accidental leakage of secret data. We discuss additional options to secure extraction further in § 8.

### v4

@@ Abstract.
Capabilities, constant-time code, data-oblivious code, hardware implementation, side-channel analysis, transient-execution attacks
journalyear: 2025copyright: ccconference: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security (CCS ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3719027.3765169isbn: 979-8-4007-1525-9/2025/10
@@ 1. Introduction
Weak memory-safety and side-channel leakage are two significant security challenges for security-critical software, such as cryptographic libraries. Lack of memory-safety in programming languages such as C and C++, is one of the oldest, most persistent problems in computer security. The urgency of addressing the impact of memory-unsafe code at scale has grown under recent regulatory scrutiny (ONCD, 2024) leading cybersecurity authorities like US CISA (CISA, 2023) and UK NCSC (NCSC, 2024) to advocate for memory-safe languages like Rust (Klabnik and Nichols, 2018), and memory-safe hardware, such as Capability Hardware Enhanced RISC Instructions (CHERI) (Watson et al., 2019) as potential ways to eliminate entire classes of memory-safety vulnerabilities.
@@
Memory-safe hardware, such as CHERI, effectively addresses memory safety, but does not inherently protect against side-channel attacks. Previous work to harden CHERI against transient execution attacks (Fuchs et al., 2024) does not protect against non-transient side-channel attacks. On the other hand, inherent resistance against side channels would also be effective against transient execution attacks that employ such side channels for inferring their effects on microarchitectural properties like cache state. In principle, CHERI-capable hardware could be complemented with existing data-oblivious computation solutions like OISA (Yu et al., 2019a) or BliMe (ElAtali et al., 2024b) that protect against side channels. However, these approaches introduce considerable performance overhead because they require the addition of hardware-managed tag bits to memory.
@@
Source code artifacts for the BLACKOUT hardware and software stack are available at https://github.com/blindedcapabilities.
@@ 2.1. The CHERI capability architecture
➀ 1-bit Validity tag
➁Permissions
➂Object type
➃Bounds
➄Baseline architecture address
@@
CheriBSD (Davis et al., 2019) is a modified version of the open-source FreeBSD operating system, designed to support CHERI-RISC-V both in emulation and on hardware. It is a fully functional OS prototype, demonstrating how CHERI support can be integrated into a conventional OS design. The CheriBSD kernel and userspace can be built in “pure-capabilityCHERI C/C++ which means all conventional memory pointers are replaced by corresponding capabilities by the CHERI-LLVM compiler. In pure-capability mode, compiler-generated code derives bounded capabilities for automatic (stack-allocated) variables from the csp. For global variables, the run-time linker populates into a capability-aware global offset table (GOT), also referred to as the captable. Capabilities for heap-allocated data is handled by Cornucopia which is integrated into CheriBSD via a shim layer, known as the malloc revocation shim (MRS), which sits on top of the underlying BSD libc’s memory allocator. Consequently, CheriBSD enforces both spatial and temporal memory safety for CHERI-enabled software.
2.2. Side-Channel Leakage
@@
A prominent example of timing side channels is cache timing, which leaks information by measuring cache access latency. By comparing the latency to a predetermined threshold, software can determine whether an address is already cached by the CPU. Prime+Probe (Osvik et al., 2006), FLUSH+RELOAD (Yarom and Falkner, 2014) and Flush+Flush (Gruss et al., 2016) are example side-channel attacks using different techniques to make this timing difference dependent on secret data and extract it from the system. Transient attacks like Spectre (Kocher et al., 2019) and Meltdown (Lipp et al., 2018), expose new ways to access (architecturally inaccessible) secrets and leak them using side channels like cache timing. Speculative taint-tracking techniques (Yu et al., 2019b), such as SpectreGuard (Fustos et al., 2019), ConTExT (Schwarz et al., 2020), and ProSpeCT (Daniel et al., 2023), mitigate Spectre-style attacks by delaying instructions dependent on speculatively loaded secrets, but mitigating side-channels in general falls on the developer.
Preventing side-channel leakage is difficult for software developers. Software is often written in high-level languages that abstract away the individual CPU instructions operating on secret data, leaving decisions to the underlying compilers and software stack. Even lower-level languages like C do not provide developers with sufficient controls to prevent side-channel leakage, as the compiler can introduce side channels in the generated assembly. Furthermore, even with seemingly side-channel-free assembly, hardware optimizations transparent to software (and often undocumented), such as speculation and out-of-order execution, can inadvertently leak secrets, as described above with the Spectre vulnerabilities. The challenge of preventing side-channel leakage is therefore present at every level of the software and hardware stack, and must be solved through hardware-software co-design.
@@ 2.3. Data-Oblivious Instruction Set Architectures
Data-oblivious computation refers to algorithms and software that process data in a manner where their control-flow and memory access patterns are not affected by their input data. Constant-time code, which hardens software against side-channels, relies on data-oblivious algorithms and that the CPU cycles consumed by individual hardware instructions are independent of their operands’ values111“Constant-time code” is a misnomer, as even side-channel-resistant code may still exhibit variable latency if the latency differences are not due to the (secret) data values. Common sources of such latencies include instruction-level latencies due to frequency scaling, or unpredictable contention of data buses for loads and stores.. Historically, developers of security-critical code sensitive to side-channels, e.g., cryptographic implementations, have relied on obscure information on instruction timing differences obtained through extensive experiments on different ISAs (Pornin, 2018a). In recent years, major processor manufacturers have began to document instructions with data-operand independent timing (DIT) (Intel, 2023a) and incorporate DIT modes into their designs (Intel, 2023b; lowRISC, 2025).
The challenge of ensuring the correctness of data-oblivious software implementations, however, remains. Verifying whether a program written in a high-level language is data-oblivious has a number of challenges (§ 9). Chief among them is that data-obliviousness can only be defined at a machine-code level. This observation has motivated the creation of data-oblivious ISAs that, similar to data operand independent timing modes, provide modes of operation for CPUs to ensure that instructions operating on sensitive data do so in a data-oblivious manner. Previous works use hardware-enforced taint tracking using “blindedness tags” (ElAtali et al., 2024a) and dedicated memory partitions (Yu et al., 2019a). These approaches result in best-case performance overheads from 8% to 35% (worst case involving an order of magnitude slowdown). Memory tagging exhibits poor performance as the number of tag bits increase. Schemes such as the Arm architecture’s Memory Tagging Extension (MTE), used for memory-safety sanitation, limit the number of tag bits to 4. MTE further mitigates performance impact by allowing tag checks to occur asynchronously. But asynchronous tag checks are not practical for maintaining data-obliviousness as data-oblivious ISAs must be able to prevent data accesses before confidentiality is violated.
@@ 3. System and Adversary Model
We assume the same system model as in CHERI. We trust the OS to prevent memory containing sensitive information from being exposed to other processes after a process exits, either normally or as a result of CHERI exceptions. We assume memory-safety properties provided by CHERI, which prevent adversaries from tampering with the program’s control flow or directly inferring memory contents beyond that which are exposed during normal program operation. We also assume a DIT (lowRISC, 2025) mode is available.
We assume the same adversary model as in CHERI. In addition, we assume the adversary has the ability to observe side channels during a program’s execution, possibly from another process running on the system simultaneously. In this work, we consider side-channel threats that can be mitigated by data-oblivious software, including timing side channels, such as cache timing (Osvik et al., 2006; Yarom and Falkner, 2014; Gruss et al., 2016; Chiang and Li, 2025), instruction timing (Pornin, 2018b) (with DIT) and port-contention timing (Aldaya et al., 2019). We also consider transient execution attacks that leak information loaded by mis-speculated instructions within the same address space, such as Spectre (Kocher et al., 2019; Maisuradze and Rossow, 2018; Koruyeh et al., 2018; Horn, 2018), in scope, even if mis-training can be done across address spaces (Canella et al., 2019a). Attacks with transient instructions that can load data across address spaces, such as Meltdown and microarchitectural data sampling (including load-store-buffers) (Canella et al., 2019b), and transient capability forgery, such as Meltdown-CF (Fuchs et al., 2024), are out of scope, but can be prevented by using Capability Speculation Contracts (CSCs) (Fuchs et al., 2024). Attacks that require physical access to the system, e.g., to measure power consumption, or those through direct memory access (DMA) peripherals are out of scope.
@@ 4.1. Goals and Primary Challenges
To address these challenges, we propose enhancing the existing CHERI capability model with blinded capabilities which ensure data accessed through them is operated on exclusively in a data-oblivious manner. Memory managed by blinded capabilities is referred to as blinded memory. This avoids the need to propagate blindedness tags to memory thus overcoming the drawbacks of simply integrating an existing data-oblivious ISA with CHERI (ElAtali et al., 2024b; Yu et al., 2019a).
@@ 5.1. Software Architecture
Figure 2 shows a high-level overview of the blinded capability software stack. Developers can indicate to the compiler that a variable is sensitive by annotating its declaration with the blinded attribute ➀. We refer to such variables as blinded variables. The compiler will then ensure that any accesses to blinded variables are permitted only through blinded capabilities, guaranteeing that the values of these variables will always be blinded when loaded into registers. We refer to registers holding blinded data as blinded registers. Blinded memory can be dynamically allocated using the blinded allocator application programming interface (API) ➁ which returns a blinded capability. Similarly, the compiler enforces that references to blinded memory cannot be assigned to non-blinded variables. Inside the compiler, we modify the Clang front-end to emit blinded attributes to the LLVM intermediate representation (IR) of the compiled program ➃. These are used by the blinded capability analysis ➄ and instrumentation passes ➅ to produce a blinded capability-instrumented application binary. Global variables annotated with blinded attribute are reserved from a separate blinded data section by the static linker.
At run-time, the blinded-capability-enhanced application ➆ requires platform support in the form of the blinded allocator ➇ and an OS with a CHERI-capable kernel that has been altered to ensure its internal handling of capabilities does not trigger blinded capability violations. C startup code added by the compiler to ➆ is responsible for initializing capabilities pointing to data in the blinded data section as blinded capabilities. System software not making use of blinded capabilities does not require modification (apart from their adaptation to CHERI). We discuss the compiler enhancements and operating system changes in detail in § 6.2.
1#define __blinded [[clang::annotate_type("blinded")]] 2 3 __attribute__ ((blinded)) 4int data_oblivious_select(bool cond, int x, int y) { 5 6 bool __blinded c; // c declared blinded and 7 // accessed via blinded capability 8 9 int res; // res not declared blinded 10 // but blindedness is inferred 11 12 13 c = cond; // Uses store via blinded capability 14 // (argument already in register) 15 16 17 res = (x * c) + (y * (!c)); // HW propagates 18 // blindedness to res 19 20 21 return res; 22} Listing 1: Data oblivious conditional select function which returns one of the arguments x or y based on the value of cond demonstrating the inference capabilities of the blinded capability-enhanced compiler. ❸ Compiler infers res blinded from ❺❹ Compiler knows c is already blinded based on declaration❺ Compiler infers res is blinded from this assignment
@@
Table 1. Blindedness bit propagation and side-channel prevention rules enforced by BLACKOUT hardware. x represents a “don’t care” condition, which indicates the actual signal value or values have no impact on the decision outcome.
Instruction Blinded Blindedness Decision Result
capability blindedness
Arithmetic / - Op1 Op2
Logic $a$ $b$ Propagate $a\vee b$
Branching in Addr Reg Addr Reg Condition Ops
no 0 0 Allow -
no 1 x Fault -
no x 1 Fault -
yes x x Fault -
Load in Addr Reg Addr Reg
no 0 Propagate 0
yes 0 Propagate 1
x 1 Fault -
Store in Addr Reg Addr Reg Data Reg
no 0 0 Allow -
no 0 1 Fault -
yes 0 x Allow -
x 1 x Fault -
@@ 5.2. Hardware architecture
Figure 3 shows the high-level overview of the hardware changes needed to support blinded capabilities (shown in dark gray) applied to an out-of-order RISC-V core, such as MIT’s RISCY-OO (Zhang, 2025) or Bluespec’s Toooba (Nikhil, 2025) processor intellectual property (IP). Apart from the changes introduced by CHERI (shown in light gray) the majority of the changes necessary in the CPU are to facilitate in-CPU taint tracking. In the register file, the general-purpose capability registers are extended to hold an additional blindedness bit ➀ which tracks whether the register is blinded. The taint-propagation hardware logic ➁ will in turn ensure that any data derived from these blinded registers, e.g., through arithmetic operations, will also “be blinded”, i.e., it sets the blindedness bit of the destination register.
All blinded data is bound by a set of invariants (I1, I2, I3, I4 and I5), stemming from the need to enforce exclusive access to blinded memory (§ 4.2), and side-channel prevention.
@@
I1 is enforced by hardware through additional blinded store and load checks ➂ in the memory execution pipeline which is responsible for the execution of any load and store instructions, whereas I4 and I5 are enforced by side-channel prevention logic integrated into any pipeline which is responsible for instructions with blinded operands ➃. Any attempt by software to violate I1, I4 and I5 will result in a hardware fault. Table 1 shows the full blindedness bit propagation and side-channel prevention rules enforced by BLACKOUT.
@@
Data-oblivious computation ensures that control flow and memory accesses don’t depend on secret data. Outputs, however, may be non-secret (e.g., decrypted data meant for users). In practice, such output data cannot be left blinded indefinitely as some result of data-oblivious computation must eventually be extracted. In BLACKOUT, results can be marked non-secret by issuing non-blinded capabilities for result buffers. This reveals the final, non-secret result of blinded computation by relaxing invariant I3 in a controlled way to unblind the result. For example, since I3 is (in part) enforced by the blinded allocator, the allocator can expose a separate API for obtaining blinded results that, in contrast to, regular blinded memory returns two capabilities for a dedicated area for blinded results: one which is blinded, and used to write to the area, and the other non-blinded, but only usable for reading the result. However, the developer must use the blinded capability to the result area carefully to prevent accidental leakage of secret data. We discuss additional options to secure extraction further in § 8.